# The integral form of Gauss’ law for a generic case of the electrodynamics

### v1

@@ Abstract
We have analytically examined the integral form of Gauss’ law for arbitrarily moving charges inside and outside an arbitrarily moving and deforming Gaussian surface with the Maxwell equations under consideration. We have obtained a simple looking differential equation $\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{I^{(s)}_{\text{in}}(t)}{\epsilon_{0}}$ for time-dependent Gauss’ flux-integral. We have explicitly calculated the time-derivative of the flux-integral for the electric field passing through the Gaussian surface for such a generic non-static case and have found no changes in the form of Gauss’ law.
pacs:
01.55.+b General physics, 03.50.De Classical electromagnetism, Maxwell equations, 01.40.Fk Research in physics education
Classical electrodynamics is a well-established branch of physics and engineering since its one and half centuries old (1865 AD) formulation with the Maxwell equations Born . While the integral form of Gauss’ law (1835 AD) is a pillar for electrostatics, the differential form of Gauss’ law is a pillar for the electrodynamics because it is one of the four Maxwell equations Born . This article, however, is related to the formulation of the classical electrodynamics because we are going to examine the validity of the integral form of Gauss’ law for a generic time-dependent case with the Maxwell equations under consideration.
@@
in the standard textbooks Jackson ; Griffiths for a static case (i.e. for $\frac{\partial\rho}{\partial t}=0$ and $\frac{\partial\vec{J}}{\partial t}=0$) of the total change $q_{\text{in}}$ inside a (closed) Gaussian surface $s$ where a point is denoted by $\vec{r}^{\prime}$. Laws of electrostatics and magnetostatics such as Biot-Savart law and Coulomb’s law, respectively, were already generalized for the electrodynamics with the time-dependence under consideration Griffiths2 . The differential form of Gauss’ law, i.e. Eqn. (1), on the other hand, is valid for a generic case of electrodynamics Born . Thus, a natural question comes - whether the differential form of Gauss’ law, i.e. Eqn. (5), is also valid for a generic non-static (dynamic) case of the charges moving arbitrarily inside and outside an arbitrarily moving and deforming Gaussian surface. A special case of such a study with a fixed amount of charges inside the moving and deforming surface, however, was discussed before Rothwell . Two even more special cases of such a study were also separately discussed (i) for a single charge uniformly moving and (ii) for an accelerated radiating charge, inside a static Gaussian surface under the consideration of the retarded scalar potential Garcia .
@@
The left-hand side of Eqn. (5) represents the flux of the electric field passing through the Gaussian surface $s$. For a non-static case as depicted in figure 1, the flux of the electric field passing through a moving Gaussian surface $s(t)$ can be given by the time-dependent integral $\oint_{s(t)}\vec{E}(\vec{r}^{\prime},t)\cdot\text{d}\vec{s}(t)$. This flux-integral may vary with time if the Gaussian surface element $\text{d}\vec{s}(t)$ at position $\vec{r}^{\prime}$ and time $t$ moves with a velocity $\vec{v}^{\prime}(\vec{r}^{\prime},t)$ and the (point) charges $\{q_{i}\}$ ($i=0,1,2,3,...,n$) both inside and outside $s(t)$, as depicted in figure 1, move with the velocities $\{\vec{v}_{i}(\vec{r}_{i},t)\}$. By notation, we have $\vec{r}=\vec{r}^{\prime}$ for $\vec{r}$ on the surface $s(t)$. Otherwise, $\vec{r}$ denotes a point inside and outside $s(t)$. The time-derivative of the flux-integral can be obtained by following the procedure described in Ref.Jackson1 ; Zangwill as
@@
Now, using the Maxwell equations, especially Eqns. (1) and (4), we recast Eqn. (The integral form of Gauss’ law for a generic case of the electrodynamics) as
@@
where $\triangle q_{\text{in}}^{(s)}(t,t_{0})=\int_{t_{0}}^{t}I^{(s)}_{\text{in}}(t^{\prime})\text{d}t^{\prime}$ is the net charge accumulated inside the Gaussian surface $s(t)$ through the entire moving surface $s(t^{\prime})$ in the course of time for $t_{0}<t^{\prime}<t$. Let us now suitably choose the reference time $t_{0}$ such that both the charges and the Gaussian surface were static from $t=-\infty$ to $t=t_{0}$ with the total amount of charge $q_{in}^{(0)}$ in it. Thus, the flux-integral $\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})$ follows Gauss’ law $\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})=\frac{q_{\text{in}}^{(0)}}{\epsilon_{0}}$ as mentioned in Eqn. (5) for the electrostatics Jackson ; Griffiths . Thus, we recast Eqn. (10) as
$\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{q_{\text{in}}(t)}{\epsilon_{0}}$ (11)
where $q_{\text{in}}(t)=q_{\text{in}}^{(0)}+\triangle q_{\text{in}}^{(s)}(t,t_{0})$ is the net charge inside the Gaussian surface $s(t)$ at any instant of time $t$ for $t>t_{0}$. Eqn. (11) is the desired result for a generic non-static case of charges moving inside and outside an arbitrarily moving and deforming Gaussian surface. It is clear from Eqn. (11) and Eqn. (5) that the integral form of Gauss’ law remains unaltered, i.e. robust, even for a non-static case with a moving and deforming Gaussian surface. The total amount of charges inside the moving surface $s(t^{\prime})$ (for $t_{0}<t^{\prime}<t$) was not necessarily considered to be conserved for the derivation of Eqn. (11). The total charge inside the moving surface would have been conserved if we had considered $I^{(s)}_{\text{in}}(t)=0$ for $t_{0}<t^{\prime}<t$.
Let us now consider a special case of the conservation of the total charge inside the moving Gaussian surface. It is interesting to note from Eqn. (11) that, if the charges arbitrarily move inside the arbitrarily moving and deforming surface $s(t^{\prime})$ and no current passes through it for $t_{0}<t^{\prime}<t$, then we have $\triangle q_{\text{in}}^{(s)}(t,t_{0})=0$ for $t_{0}<t^{\prime}<t$. In such a special case, Eqn. (11) takes the form
@@
It should be mentioned that the flux of the electric field through a part of the moving Gaussian surface $s(t^{\prime})$ though changes in time for $t_{0}<t^{\prime}<t$, the flux through whole the Gaussian surface, of course, does not change at all in such a special case. This special case [^1] The case was made special in Ref.Rothwell for the use of the continuity equation for the charges inside the moving surface. was studied in Ref.Rothwell without an explicit derivation. Eqn. (12) is the same as that obtained in Ref.Rothwell . Eqn. (12) is also compatible with the previous study for the even more special cases of the uniform (relativistic) motion of a charge inside a static Gaussian surface and the accelerated motion of a radiating charge inside a static Gaussian surface Garcia .
To conclude, we have obtained a simple looking differential equation for time-dependent Gauss’ flux-integral for charges arbitrarily moving inside and outside an arbitrarily moving and deforming Gaussian surface. There-from we have explicitly calculated the time-dependent flux-integral for the electric field passing through the Gaussian surface and have found robustness in the original form of Gauss’ law.
Our calculations are based on the Maxwell equations which don’t need any relativistic correction. Hence our final result (i.e. Eqn. (11)) on the time-dependent form of Gauss’ law is essentially correct from the relativistic point of view.
The charges inside and outside the arbitrarily moving and deforming Gaussian surface are considered to be moving arbitrarily. While the uniformly moving charges result in no radiation, the accelerating charges inside and outside the Gaussian surface result in electromagnetic radiation. This radiation does not change the form of Gauss’ law as clear from Eqn. (11).
Finally, it should be mentioned that the differential equation for Gauss’ flux-integral (Eqn. (9)) obtained by us, is novel, simple looking, and important for determining time-dependent Gauss’ flux-integral. The time-dependent form of Gauss’ flux-integral (Gauss’ law), i.e. Eqn. (11), obtained by us, is also absolutely new. Our explicit derivation of Gauss’ flux-integral for a generic non-static case, however, can be directly useful even in an undergraduate or postgraduate physics class on electrodynamics for a better understanding of Gauss’ law. Gauss’ flux-integral obtained by us would be applicable for arbitrary charge and current distributions with radiating charges and non-conserved total charge in a region.
@@ Acknowledgement
S. Biswas acknowledges the partial financial support of the SERB (now ANRF), DST, Govt. of India under the EMEQ Scheme [No. EEQ/2023/000788]. We thank Mr. Bikram Keshari Behera (UoH, India) for helping us drawing figure 1.

### v2

# Revisiting the integral form of Gauss’ law for a generic case of electrodynamics

@@ Abstract
We have analytically examined the integral form of Gauss’ law for arbitrarily moving charges inside and outside an arbitrarily expanding (or contracting) and deforming Gaussian surface with explicit use of the motion of the Gaussian surface and the Maxwell equations under consideration. We have obtained a simple-looking differential equation $\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{I^{(s)}_{\text{in}}(t)}{\epsilon_{0}}$ for time-dependent Gauss’ flux-integral. We have explicitly calculated time-dependent Gauss’ flux-integral for the moving surface and have found no changes in the original result which was obtained without explicit use of the motion. Our explicit derivation of Gauss’ flux-integral for the moving surface can be directly useful in an undergraduate/postgraduate physics class for a better understanding of Gauss’ law for electrodynamics.
I Introduction
Classical electrodynamics is a well-established branch of physics and engineering since its one and half centuries old (1865 AD) formulation with the Maxwell equations Born . While the integral form of Gauss’ law (1835 AD) is a pillar for electrostatics, the differential form of Gauss’ law is a pillar for electrodynamics because it is one of the four Maxwell equations Born . Classical electrodynamics is a very common undergraduate/postgraduate course for both the physics and the engineering students. The classical electrodynamics with exotic or semi-classical properties of the optical, dielectric, magnetic, conducting, etc media, however, is still a hunting ground for theoretical and experimental research Garcia-Vidal ; Basov ; Lapine ; Rotter ; Garcia-Vidal2 . The integral form of Gauss’ law in electrodynamics is found in the textbooks Rothwell2 . However, students often ask questions on the integral form of Gauss’ law in electrodynamics especially when the charges inside the Gaussian surface move arbitrarily and become radiative. Questions may also come if the Gaussian surface moves further. In this article, we would like to examine the validity of the integral form of Gauss’ law for a generic time-dependent case of charges moving arbitrarily inside and outside a moving (expanding/contracting and deforming) Gaussian surface with the Maxwell equations under consideration.
@@ I Introduction
in the standard textbooks Jackson ; Griffiths for a static case (i.e. for $\frac{\partial\rho}{\partial t}=0$ and $\frac{\partial\vec{J}}{\partial t}=0$) of the total change $q_{\text{in}}$ inside a static (closed) Gaussian surface $s$ where a point is denoted by $\vec{r}^{\prime}$. However, if we integrate Eqn. (1) with respect to $\text{d}^{3}\vec{r}$ in the volume $V(t)$ and apply Gauss’s divergence theorem on the left-hand side, we can easily get
$\oint_{s(t)}\vec{E}(\vec{r}^{\prime},t)\cdot\text{d}\vec{s}(t)=\frac{q_{\text{in}}(t)}{\epsilon_{0}}$ (6)
where $q_{\text{in}}(t)=\int_{V(t)}\rho(\vec{r},t)\text{d}^{3}\vec{r}$ is the total (arbitrarily moving) charge inside the volume $V(t)$ at any instant of time $t$ and $s(t)$ represents the moving Gaussian surface which encloses the volume $V(t)$ in the free space as described in figure 1. It appears that the form of Eqn. (6) for a generic non-static case is essentially the same as that of Eqn. (5) for a static case. Let us call Eqn. (6) as the original integral form of Gauss’s law for electrodynamics. This time-dependent form is also available in textbooks Rothwell2 . However, the time $t$ has been used as a parameter in both the $s(t)$ and the $V(t)$ while deriving Eqn. (6) from Eqn. (5). Such a parametrization does not explicitly take neither the motion of the surface $s(t)$ nor the motion of the charges inside and outside $s(t)$ into account. Thus a doubt remains whether Eqn. (6) is true for a generic non-static case of both the Gaussian surface $s(t)$ and the charges inside and outside the Gaussian surface. Hence we are motivated to examine the validity of Eqn. (6) with the explicit use of motion of both the Gaussian surface $s(t)$ and the charges inside and outside $s(t)$.
@@
Laws of electrostatics and magnetostatics such as Biot-Savart law and Coulomb’s law, respectively, were already generalized for electrodynamics with the time-dependence under consideration Griffiths2 . The differential form of Gauss’ law, i.e. Eqn. (1), on the other hand, is valid for a generic case of electrodynamics Born . Thus, a natural question comes - whether the differential form of Gauss’ law, i.e. Eqn. (5), is also valid for a generic non-static (dynamic) case of the charges moving arbitrarily inside and outside an arbitrarily moving Gaussian surface. A special case of such a study with a fixed amount of charges inside the moving surface, however, was discussed before Rothwell . Two even more special cases of such a study were also separately discussed (i) for a single charge uniformly moving and (ii) for an accelerated radiating charge, inside a static Gaussian surface under the consideration of the retarded scalar potential Garcia .
Our calculation in this article starts with the time-derivative of the flux-integral under the consideration of the Maxwell equations as described in Refs.Jackson1 ; Zangwill for magnetic field passing through an unclosed moving surface. There-from we obtain a simple-looking differential equation $\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{I^{(s)}_{\text{in}}(t)}{\epsilon_{0}}$ for time-dependent Gauss’ flux-integral for the electric field passing through a moving (expanding/contracting and deforming) Gaussian surface where-in the total amount of charges is a non-conserved quantity. Then we integrate this equation and obtain time-dependent Guass’ flux integral for the generic case of electrodynamics. Finally, we conclude.
II Gauss’ flux-integral in electrodynamics
@@
The left-hand side of Eqn. (5) represents the flux of the electric field passing through the Gaussian surface $s$. For a non-static case as depicted in figure 1, the flux of the electric field passing through a moving (expanding/contracting and deforming) Gaussian surface $s(t)$ can be given by the time-dependent integral $\oint_{s(t)}\vec{E}(\vec{r}^{\prime},t)\cdot\text{d}\vec{s}(t)$. This flux-integral may vary with time as the surface element $\text{d}\vec{s}(t)$ at the position $\vec{r}^{\prime}$ at time $t$ moves with a velocity $\vec{v}^{\prime}(\vec{r}^{\prime},t)$ and the (point) charges $\{q_{i}\}$ ($i=0,1,2,3,...,n$) both inside and outside $s(t)$, as depicted in figure 1, move with the velocities $\{\vec{v}_{i}(\vec{r}_{i},t)\}$. By notation, we have $\vec{r}=\vec{r}^{\prime}$ for $\vec{r}$ on the surface $s(t)$. Otherwise, $\vec{r}$ denotes a point inside and outside $s(t)$. The time-derivative of the flux-integral can be obtained by following the procedure described in Refs.Jackson1 ; Zangwill as
@@
While the 2nd term on the right-hand side of Eqn. (II) is coming from the expansion (or contraction) of the Gaussian surface $s(t)$, the 3rd term on the same side is coming from its deformation Jackson1 ; Zangwill . Now, using the Maxwell equations, especially Eqns. (1) and (4), we recast Eqn. (II) as
@@
where $\triangle q_{\text{in}}^{(s)}(t,t_{0})=\int_{t_{0}}^{t}I^{(s)}_{\text{in}}(t^{\prime})\text{d}t^{\prime}$ is the net charge accumulated inside the Gaussian surface $s(t)$ through the entire moving surface $s(t^{\prime})$ in the course of time for $t_{0}<t^{\prime}<t$. Let us now suitably choose the reference time $t_{0}$ such that both the charges and the Gaussian surface were static from $t=-\infty$ to $t=t_{0}$ with the total amount of charge $q_{in}^{(0)}$ in it. Thus, the flux-integral $\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})$ follows Gauss’ law $\oint_{s(t_{0})}\vec{E}\cdot\text{d}\vec{s}(t_{0})=\frac{q_{\text{in}}^{(0)}}{\epsilon_{0}}$ as mentioned in Eqn. (5) for electrostatics Jackson ; Griffiths . Thus, we recast Eqn. (11) as
$\oint_{s(t)}\vec{E}(\vec{r}^{\prime},t)\cdot\text{d}\vec{s}(t)=\frac{q_{\text{in}}(t)}{\epsilon_{0}}$ (12)
where $q_{\text{in}}(t)=q_{\text{in}}^{(0)}+\triangle q_{\text{in}}^{(s)}(t,t_{0})$ is the net charge inside the Gaussian surface $s(t)$ at any instant of time $t$ for $t>t_{0}$. Eqn. (12) is the desired result for a generic non-static case of charges moving inside and outside an arbitrarily moving (expanding/contracting and deforming) Gaussian surface. By comparing Eqns. (12) and (6), we can safely say that it does not require any change in the original Gauss’ law which was obtained without explicit use of the motion. The total amount of charges inside the moving surface $s(t^{\prime})$ (for $t_{0}<t^{\prime}<t$) has not been necessarily considered to be conserved for the derivation of Eqn. (12). The total charge inside the moving surface would have been conserved if we had considered $I^{(s)}_{\text{in}}(t)=0$ for $t_{0}<t^{\prime}<t$.
Let us now consider a special case of the conservation of the total charge inside the moving Gaussian surface. It is interesting to note from Eqn. (12) that, if the charges arbitrarily move inside the arbitrarily moving surface $s(t^{\prime})$ and no current passes through it for $t_{0}<t^{\prime}<t$, then we have $\triangle q_{\text{in}}^{(s)}(t,t_{0})=0$ for $t_{0}<t^{\prime}<t$. In such a special case, Eqn. (12) takes the form
@@
It should be mentioned that the flux of the electric field through a part of the moving Gaussian surface $s(t^{\prime})$ though changes in time for $t_{0}<t^{\prime}<t$, the flux through the whole Gaussian surface, of course, does not change at all in such a special case. This special case [^1] The case was made special in Ref.Rothwell for the use of the continuity equation for the charges inside the moving surface. was studied in Ref.Rothwell without an explicit derivation. Eqn. (13) is the same as that obtained in Ref.Rothwell . Eqn. (13) is also compatible with the previous study for the even more special cases of the uniform (relativistic) motion of a charge inside a static Gaussian surface and the accelerated motion of a radiating charge inside a static Gaussian surface Garcia .
III Conclusion
To conclude, we have analytically examined the integral form of Gauss’ law for arbitrarily moving charges inside and outside an arbitrarily expanding (or contracting) and deforming Gaussian surface with explicit use of the motion of the Gaussian surface and the Maxwell equations under consideration. We have obtained a differential equation $\frac{\text{d}}{\text{d}\text{t}}\oint_{s(t)}\vec{E}\cdot\text{d}\vec{s}(t)=\frac{I^{(s)}_{\text{in}}(t)}{\epsilon_{0}}$ for time-dependent Gauss’ flux-integral. There-from, we have explicitly calculated time-dependent Gauss’ flux-integral for the moving surface and have found no changes in the original result which was obtained without explicit use of the motion.
Our calculations are based on the Maxwell equations which don’t need any relativistic correction. Hence the final result (i.e. Eqn. (12)) on the time-dependent form of Gauss’ law is essentially correct from the relativistic point of view.
The charges inside and outside the arbitrarily moving Gaussian surface are considered to be moving arbitrarily. While the uniformly moving charges result in no radiation, the accelerating charges inside and outside the Gaussian surface result in electromagnetic radiation. This radiation does not change the form of Gauss’ law as clear from Eqn. (12).
Finally, it should be mentioned that the differential equation for Gauss’ flux-integral (Eqn. (10)) obtained by us, is novel, simple-looking, and important for determining time-dependent Gauss’ flux-integral. The time-dependent form of Gauss’ flux-integral (Gauss’ law), i.e. Eqn. (12), obtained by us although distinguished with the form $q_{\text{in}}(t)=q_{\text{in}}^{(0)}+\triangle q_{\text{in}}^{(s)}(t,t_{0})$, is essentially not a new result because it exactly matches with (Eqn. (6)) with the form $q_{\text{in}}(t)=\int_{V(t)}\rho(\vec{r},t)\text{d}^{3}\vec{r}$. Our explicit derivation of Gauss’ flux-integral for the moving surface can, however, be directly useful in an undergraduate/postgraduate physics class for a better understanding of Gauss’ law for electrodynamics.
@@ Acknowledgement
S. Biswas acknowledges the partial financial support of the SERB (now ANRF), DST, Govt. of India under the EMEQ Scheme [No. EEQ/2023/000788]. We thank Mr. Bikram Keshari Behera (UoH, India) for helping us draw figure 1. Useful discussions with Dr. Abhiram Soori (UoH, India), Dr. N. Sri Ram Gopal (UoH, India), Prof. V. Subrahmanyam (UoH, India), and Prof. Koushik Ray (IACS, India) are gratefully acknowledged.